OpenStudy (anonymous):

Joe is going to travel to Dallas. He lives 320 miles away and can average 69 miles per hour. Use an equation to represent the distance D from Dallas in terms of h, the number of hours Joe has traveled.

OpenStudy (anonymous):

D = rt

OpenStudy (anonymous):

320=69t

OpenStudy (anonymous):

do you know what to do next

OpenStudy (anonymous):

about 5 hours

OpenStudy (anonymous):

Divide

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a. D = 69h c. D = 320 + 69h b. D = 320 - 69h d. D = 320h + 69

OpenStudy (anonymous):

Those are some of the multiple choice answers

OpenStudy (anonymous):

id say the 2nd one